# Skill-aware Mutual Information Optimisation for Generalisation in Reinforcement Learning

### v1

@@ Appendix B Proof for Lemma 2
Step 1. According to Lemma 1, we have ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}\leq{\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}$ (shown in Figure 3).
Step 2. Let us prove SaNCE is a $K$-sample SaNCE estimator and upper bounded by ${\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}$. Because $\frac{f_{\psi}(c,\pi_{c},\tau_{c}^{+})}{f_{\psi}(c,\pi_{c},\tau_{c}^{+})+\sum_{k=2}^{K}f_{\psi}(c,\pi_{c},\tau_{c,k}^{-})}\leq 1$ [Mu et al., 2022], so we have:
$\quad{\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$
@@
Hence, we have ${\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}$ (similar to Equation 6).
Step 3. With the definition of $K^{*}$, we can prove ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$ with the same sample size $K$. In task $e_{1}$ with context embedding $c_{1}$, SaNCE obtains positive and negative samples from the current task $e_{1}$. Hence the variable $c=c_{1}$ is constant, we have:
$\quad{\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$
@@
The required sample size $K^{*}_{\text{SaNCE}}=|c_{1}|\cdot|\pi|\cdot M=|\pi|\cdot M$. With $K\to K^{*}_{\text{SaNCE}}$, ${\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\approx{\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}$. Correspondingly, for InfoNCE, in the current task $e_{1}$ with context embedding $c_{1}$, positive samples are trajectories $\tau_{1}$ generated after executing the skill $\pi_{1}$ in task $e_{1}$, and negative samples are trajectories $\{\tau_{c_{2}}^{-},...\}$ from other tasks $\{e_{2},...\}$. Under the definition of $K^{*}$, we have:
$\quad{\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$
@@
where $K^{*}_{\text{InfoNCE}}=|c|\cdot|\pi|\cdot M\approx|c|\cdot K^{*}_{\text{SaNCE}}$. In real-world robotic control tasks, the sample space size significantly increases due to multiple environmental features $e=\{e^{0},e^{1},...,e^{N}\}$. $|c|$ is the number of different tasks and increases exponentially due to the permutations and combinations of $N$ environmental features. When current task $e_{1}$ has context embedding $c_{1}$, the $c_{2:|c|}$ refer to the context embedding for the other tasks. With $K\to K^{*}_{\text{InfoNCE}}$, ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\approx{\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}$. Hence, during the training process, ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$ with the same sample size $K$.
Step 4. What remains is to show ${\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}\leq{\color[rgb]{0.21875,0.328125,0.57421875}I(c;\tau_{c})}$. The MI for three variables is also called interaction information. According to the definition in McGill [1954], the SaMI can be presented as:
Proposition 2
For the case of three variables, the MI could be written as ${\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}={\color[rgb]{0.21875,0.328125,0.57421875}I(c;\tau_{c})}-I(c;\tau_{c}|\pi_{c})$.
Using this proposition, one can see that in the case of three variables, interaction information quantifies how much the information shared between two variables differs from what they share if the third variable is known. Several properties of interaction information in the case of three variables have been studied in the literature. Specifically, Yeung [1991] showed that
$-\min\{I(x;\tau_{c}|\pi_{c}),I(c;\pi_{c}|\tau_{c}),I(\tau_{c};\pi_{c}|c)\}\leq I(c;\tau_{c};\pi_{c})\leq\min\{I(c;\tau_{c}),I(c;\pi_{c}),I(\tau_{c};\pi_{c})\}$ (12)
Combining Equation 12 and Proposition 2, we have ${\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}\leq{\color[rgb]{0.21875,0.328125,0.57421875}I(c;\tau_{c})}$.
@@
Putting it together. Hence, we have Lemma 2: we always have ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.78125,0.67578125,0.0390625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}\leq{\color[rgb]{0.3828125,0.33984375,0.5859375}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}\leq{\color[rgb]{0.21875,0.328125,0.57421875}I(c;\tau_{c})}$ (shown in Figure 3), while learning a skill-aware context encoder $\psi$ with SaNCE estimator. $K^{*}_{\text{SaNCE}}<<K^{*}_{\text{InfoNCE}}$ so that $I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)$ is a much tighter lower bound of the true $I_{\text{SaMI}(c;\pi_{c};\tau_{c})}$ than ${\color[rgb]{0.5625,0.66796875,0.80859375}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$.
@@ Appendix C Sample size of $I_{\text{Sa+InfoNCE}}$
We illustrate the sample size of $I_{\text{Sa+InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)$ in this section. Sa+InfoNCE plugs SaNCE into InfoNCE, and has positive samples $\tau_{c_{1}}^{+}$ from task $e_{1}$ after executing skill $\pi_{c_{1}}^{+}$, and negative samples are trajectories $\tau_{c_{1:K}}^{-}$ from executing skills $\pi_{c_{1:K}}^{-}$ in task $e_{1:K}$, respectively. Therefore, this is equivalent to observing the variable $c$, then observing the variable $\pi_{c}$, i.e., sampling from the distribution $p(\pi_{c},\tau_{c}|c)p(c)$. We have:
@@
It should be noted that such a combination will increase the size of the negative sample space, i.e., $K^{*}_{\text{Sa+InfoNCE}}=\left(\sum_{i=1}^{|c|}|\pi_{c_{i}}^{-}|+|\pi_{c_{1}}^{+}|\right)\cdot M\geq K^{*}_{\text{SaNCE}}$. Therefore, with the same number $K$ of samples, $I_{\text{Sa+InfoNCE}}$ is less precise and looser than $I_{\text{SaNCE}}$. Hence, a trade-off between sample diversity and the precision of the $K$-sample estimator is required.
@@ Appendix D Environmental setup
Figure 8: (a) Modified Panda-gym benchmarks, (b) the training tasks, the moderate test tasks and the extreme test tasks. The moderate test task setting just has a combinatorial interpolation. The extreme test task setting has unseen ranges of environmental features and is an extrapolation.
@@ D.1 Modified Panda-gym
We modified the original Pick&Place task in Panda-gym [Gallouédec et al., 2021] by setting the $z$ dimension (i.e., the desired height) of the cube’s goal position equal to 00  [^3] If $z$ is not equal to 0, Pick&Place skill is always needed to solve tasks. and maintaining the freedom  [^4] In the original Push task, the grippers are blocked to ensure the agent can only push cubes. However, this hinders the agent from learning Pick&Place skills, leading to failure when it encounters an ”unpushable” scenario. of grippers, so the agent can explore whether it is supposed to push the cube or grasp it. Skills in this benchmark are defined as:
  • Pick&Place skill: This skill specifically refers to the agent trying to use the gripper to grasp the cube, pick it up off the table and place it in the goal position. We determine the Pick&Place skill by detecting no contact points between the table and the cube, two contact points between the robot’s end effector and the cube, and the cube’s height being greater than half its width, meaning it is on the table.
  • Push skill: This skill refers to the agent giving a push force to the cube and then the cube slides to the goal position. We confirm the Push skill by detecting that the cube’s height is equal to half its width.
  • Other skills: Besides the Pick&Place and Push skills, any other different modes of behaviour are classified as other skills. For example, with the Flip skill, the agent applies an initial force to the cube, causing it to roll on the surface or leave the table. In this case, we detect that the cube’s height is greater than half its width, and the number of contact points between the cube and the table is greater than zero.
Some elements in the RL framework are defined as the following:
State space: we use feature vectors which contain cube position (3 dimensions), cube rotation (3 dimensions), cube velocity (3 dimensions), cube angular velocity (3 dimensions), end-effector position (3 dimensions), end-effector velocity (3 dimensions), gripper width (1 dimension), desired goal (3 dimensions) and achieved goal (3 dimensions). Environmental features are not included in the state.
Action space: the action space has 4 dimensions, the first three dimensions are the end-effector’s position changes and the last dimension is the gripper’s width change.
During training, we randomly select a combination of environmental features from a training set sampling combinations from sets: $\text{mass}=1.0$ and $\text{friction}\in\{0.1,1.0,5.0,10.0\}$; $\text{mass}\in\{1.0,5.0,10.0\}$ and $\text{friction}=0.1$. At test time, we evaluate each algorithm in all tasks from the moderate test task setting, in which $\text{mass}\in\{5.0,10.0\}$ and $\text{friction}\in\{1.0,5.0,10.0\}$ (shown in Figure 8(b)), and all tasks from extreme test task setting: $\text{mass}=30.0$ and $\text{friction}\in\{0.1,1.0,5.0,10.0,30.0\}$; $\text{mass}\in\{1.0,5.0,10.0,30.0\}$ and $\text{friction}=30.0$ (shown in Figure 8(b)).
@@ D.2 Modified MuJoCo
We follow exactly the same environmental settings as DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020]. For Hopper, Half-cheetah, Ant, and SlimHumanoid, we use the environments from the MuJoCo physics engine and use implementation available from Clavera et al. [2019b] and Seo et al. [2020], and scale the mass of every rigid link by scale factor $m$, and scale damping of every joint by scale factor $d$. For Crippled Ant and Crippled Half-cheetah, we use implementation available from Seo et al. [2020] and scale the mass of every rigid link by scale factor $m$, scale damping of every joint by scale factor $d$, and randomly select one leg, and make it crippled to change the dynamic transitions.
Figure 9: Six environments in modified MuJoCo benchmark.
Generalisation performance is measured in two different regimes: moderate and extreme, where the moderate draws environment parameters from a closer range to the training range, compared to the extreme. We have provided our settings for training, extreme and moderate test tasks in Table 3.
Table 3: Environment parameters used for MuJoCo benchmark.
@@
Hopper $m\in\{0.75,1.0,1.25\}$ $d\in\{0.75,1.0,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ 500
Crippled Ant $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ Crippled Leg: $\{0,1,2\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ Crippled Leg: $\{3\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ Crippled Leg: $\{3\}$ 1000
Crippled Half-cheetah $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ Crippled Leg: $\{0,1,2,3\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ Crippled Leg: $\{4,5\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ Crippled Leg: $\{1\}$ 1000

### v3

@@ Appendix B Proof for Lemma 2
Step 1. According to Lemma 1, we have ${\color[rgb]{0.51953125,0.73046875,0.2578125}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}\leq{\color[rgb]{0.40625,0.140625,0.53125}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}$ (shown in Figure 3).
Step 2. Let us prove that SaNCE is a $K$-sample SaNCE estimator and is upper bounded by ${\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}$. Since $\frac{f_{\psi}(c,\pi_{c},\tau_{c}^{+})}{f_{\psi}(c,\pi_{c},\tau_{c}^{+})+\sum_{k=2}^{K}f_{\psi}(c,\pi_{c},\tau_{c,k}^{-})}\leq 1$ [Mu et al., 2022], we have:
$\quad{\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$
@@
Thus, we obtain ${\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}$, similar to Equation (6).
Step 3. With the definition of $K^{*}$, we can prove that ${\color[rgb]{0.51953125,0.73046875,0.2578125}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$ with the same sample size $K$. In task $e_{1}$, SaNCE obtains positive and negative samples from the current task $e_{1}$. Since the variable $c=c_{1}$ is constant, we have:
$\quad{\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$
@@
The required sample size is $K^{*}_{\text{SaNCE}}=|c_{1}|\cdot|\pi|\cdot M=|\pi|\cdot M$. As $K\to K^{*}_{\text{SaNCE}}$, we have ${\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\approx{\color[rgb]{0.40625,0.140625,0.53125}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}$. Correspondingly, for InfoNCE, in the current task $e_{1}$ with context embedding $c_{1}$, positive samples are trajectories $\tau_{1}$ generated after executing the skill $\pi_{1}$ in task $e_{1}$, while negative samples are trajectories $\{\tau_{c_{2}}^{-},...\}$ from other tasks $\{e_{2},...\}$. Under the definition of $K^{*}$, we have:
$\quad{\color[rgb]{0.51953125,0.73046875,0.2578125}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$
@@
where $K^{*}_{\text{InfoNCE}}=|c|\cdot|\pi|\cdot M\approx|c|\cdot K^{*}_{\text{SaNCE}}$. In real-world robotic control tasks, the sample space size increases significantly due to multiple environmental features $e=\{e^{0},e^{1},\dots,e^{N}\}$. The number of different tasks $|c|$ grows exponentially due to the permutations and combinations of the $N$ environmental features. When the current task $e_{1}$ has context embedding $c_{1}$, the $c_{2:|c|}$ refer to the context embeddings for the other tasks. As $K\to K^{*}_{\text{InfoNCE}}$, we have ${\color[rgb]{0.51953125,0.73046875,0.2578125}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\approx{\color[rgb]{0.40625,0.140625,0.53125}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}$. Thus, during the training process, ${\color[rgb]{0.51953125,0.73046875,0.2578125}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$ with the same sample size $K$.
Step 4. According to the definition of SaMI in Equation (2), we have ${\color[rgb]{0.40625,0.140625,0.53125}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}\leq{\color[rgb]{0.265625,0.5234375,0.78125}I(c;\tau_{c})}$, as illustrated using Venn diagrams in Figure 8 (a) and (b). SaMI is formulated based on interaction information [McGill, 1954], which aims to capture the relationships among multivariate variables by quantifying the amount of information (redundancy or synergy) shared among three variables. Interaction information has been less extensively studied, partly due to its challenging interpretation from both information theory and neuroscience perspectives, as it can be either positive or negative [Li, 2022].
Figure 8: Venn diagrams illustrating (a) mutual information $I(c;\tau_{c})$, (b) interaction information $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$, and (c) the MDP graph of the context embedding $c$, skill $\pi_{c}$, and trajectory $\tau_{c}$, which represents a common-cause structure [Neuberg, 2003].
In the Meta-RL setting, the MDP (causal) graph structure, shown in Figure 8 (c), illustrates that the context embedding $c$, skill $\pi_{c}$, and trajectory $\tau_{c}$ form a common-cause structure, where $c$ acts as a shared cause influencing both the skill $\pi_{c}$ and the trajectory $\tau_{c}$. Consequently, $I(\pi_{c},\tau_{c}\mid c)<I(\pi_{c},\tau_{c})$. Therefore, $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})>0$ is guaranteed, making it interpretable in real-world robotic control tasks.
@@
Putting it together. Thus, we establish Lemma 2: we always have ${\color[rgb]{0.51953125,0.73046875,0.2578125}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}\leq{\color[rgb]{0.49609375,0.49609375,0.49609375}\log K}\leq{\color[rgb]{0.40625,0.140625,0.53125}I_{\text{SaMI}}(c;\pi_{c};\tau_{c})}\leq{\color[rgb]{0.265625,0.5234375,0.78125}I(c;\tau_{c})}$, as shown in Figure 3, while learning a skill-aware context encoder $\psi$ with the SaNCE estimator. Since $K^{*}_{\text{SaNCE}}\ll K^{*}_{\text{InfoNCE}}$, ${\color[rgb]{0.859375,0.70703125,0.15625}I_{\text{SaNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$ serves as a much tighter lower bound for the true $I_{\text{SaMI}}(c;\pi_{c};\tau_{c})$ than ${\color[rgb]{0.51953125,0.73046875,0.2578125}I_{\text{InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)}$.
@@ Appendix C Sample size of $I_{\text{Sa+InfoNCE}}$
In this section, we illustrate the sample size of $I_{\text{Sa+InfoNCE}}(c;\pi_{c};\tau_{c}|\psi,K)$. Sa+InfoNCE incorporates SaNCE into InfoNCE, using positive samples $\tau_{c_{1}}^{+}$ from task $e_{1}$ after executing skill $\pi_{c_{1}}^{+}$, and negative samples are trajectories $\tau_{c_{1:K}}^{-}$ from executing skills $\pi_{c_{1:K}}^{-}$ in tasks $e_{1:K}$, respectively. Therefore, this approach is equivalent to first observing the variable $c$ and then observing the variable $\pi_{c}$, i.e., sampling from the distribution $p(\pi_{c},\tau_{c}|c)p(c)$. We have:
@@
It should be noted that such a combination increases the size of the negative sample space, i.e., $K^{*}_{\text{Sa+InfoNCE}}=\left(\sum_{i=1}^{|c|}|\pi_{c_{i}}^{-}|+|\pi_{c_{1}}^{+}|\right)\cdot M\geq K^{*}_{\text{SaNCE}}$. The misaligned bars in Figure 4 (c) illustrate that negative sample spaces may vary across tasks. This variation arises because we define negative samples as trajectories with low returns, making the size of the negative sample space influenced by sampling randomness. With the same number $K$ of samples, $I_{\text{Sa+InfoNCE}}$ is less precise and looser than $I_{\text{SaNCE}}$. Therefore, a trade-off between sample diversity and the precision of the $K$-sample estimator is required.
@@ Appendix D Environmental setup
Figure 9: (a) Modified Panda-gym benchmarks, (b) the training tasks, moderate test tasks, and extreme test tasks. The moderate test task setting involves combinatorial interpolation, while the extreme test task setting includes unseen ranges of environmental features and represents an extrapolation.
@@ D.1 Modified Panda-gym
We modified the original Pick&Place task in Panda-gym [Gallouédec et al., 2021] by setting the $z$ dimension (i.e., the desired height) of the cube’s goal position to 00  [^3] If $z$ is not equal to 0, the Pick&Place skill is always required to solve the tasks. and maintaining the freedom of the grippers  [^4] In the original Push task, the grippers are blocked to ensure the agent can only push cubes. However, this restriction prevents the agent from learning Pick&Place skills, leading to ”unpushable” failure in Figure 1 (b)., allowing the agent to explore whether it should push or grasp the cube. Skills in this benchmark are defined as:
  • Pick&Place skill: This skill specifically refers to the agent using the gripper to grasp the cube, lift it off the table, and place it in the goal position. We determine the Pick&Place skill by detecting no contact points between the table and the cube, two contact points between the robot’s end effector and the cube, and the cube’s height being greater than half its width.
  • Push skill: This skill involves the agent moving the cube on the table to the goal position, either by dragging or sliding it. We confirm the Push skill by detecting that the cube’s height equals half its width.
  • Other skills: Any behaviour modes other than Pick&Place and Push are classified as other skills.
Some elements in the RL framework are defined as follows:
State space: We use feature vectors that contain the cube’s position (3 dimensions), cube rotation (3 dimensions), cube velocity (3 dimensions), cube angular velocity (3 dimensions), end-effector position (3 dimensions), end-effector velocity (3 dimensions), gripper width (1 dimension), desired goal (3 dimensions), and achieved goal (3 dimensions). Environmental features are not included in states.
Action space: The action space has 4 dimensions; the first three dimensions represent changes in the end-effector’s position, and the last dimension represents the change in the gripper’s width.
During training, we randomly select a combination of environmental features from a training set by sampling combinations from the following sets: $\text{mass}=1.0$ and $\text{friction}\in\{0.1,1.0,5.0,10.0\}$; $\text{mass}\in\{1.0,5.0,10.0\}$ and $\text{friction}=0.1$. At test time, we evaluate each algorithm on all tasks from the moderate test setting, where $\text{mass}\in\{5.0,10.0\}$ and $\text{friction}\in\{1.0,5.0,10.0\}$ (shown in Figure 9(b)), and on all tasks from the extreme test setting: $\text{mass}=30.0$ and $\text{friction}\in\{0.1,1.0,5.0,10.0,30.0\}$; $\text{mass}\in\{1.0,5.0,10.0,30.0\}$ and $\text{friction}=30.0$ (shown in Figure 9(b)).
@@ D.2 Modified MuJoCo
Figure 10: Ten environments in modified MuJoCo benchmark.
We extended the modified MuJoCo benchmark introduced in DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020]. In our extension, four new environments were added (Walker, Crippled Hopper, Crippled Walker, HumanoidStandup), compared to the original benchmark. Additionally, in our experiments, we used a different task set design (Table 3) than those used in the DOMINO and CaDM papers. For Hopper, Walker, Half-Cheetah, Ant, HumanoidStandup, and SlimHumanoid, we used the MuJoCo physics engine environments and implemented settings from Clavera et al. [2019b] and Seo et al. [2020], scaling the mass of every rigid link by a scale factor $m$ and the damping of every joint by a scale factor $d$. For Crippled Ant, Crippled Hopper, Crippled Walker, and Crippled Half-Cheetah, we used the implementation available from Seo et al. [2020], scaled the mass of each rigid link by a factor of $m$, scaled the damping of each joint by a factor of $d$, and randomly selected joints to be uncontrollable (i.e., masking the corresponding actions with 0). Generalisation performance is measured in two different regimes: moderate and extreme, where the moderate regime draws environmental features from a range closer to the training range compared to the extreme regime. Our settings for training, extreme, and moderate test tasks are provided in Table 3.
Table 3: Environmental features used for MuJoCo benchmark.
@@
Half-cheetah $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80,4.00\}$ $d\in\{0.20,0.40,1.60,1.80,4.00\}$ 1000
Ant $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ $m\in\{0.20,0.40,1.60,1.80,4.00\}$ $d\in\{0.20,0.40,1.60,1.80,4.00\}$ 1000
Hopper $m\in\{0.75,1.0,1.25\}$ $d\in\{0.75,1.0,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80,4.00\}$ $d\in\{0.20,0.40,1.60,1.80,4.00\}$ 1000
Crippled Hopper $m\in\{0.75,1.0,1.25\}$ $d\in\{0.75,1.0,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80,4.0\}$ $d\in\{0.20,0.40,1.60,1.80,4.0\}$ Crippled Legs $R_{1}\in\{0,1,2\}$ 1000
SlimHumanoid $m\in\{0.80,0.90,1.0,1.15,1.25\}$ $d\in\{0.80,0.90,1.0,1.15,1.25\}$ $m\in\{0.60,0.70,1.50,1.60\}$ $d\in\{0.60,0.70,1.50,1.60\}$ $m\in\{0.40,0.50,1.70,1.80\}$ $d\in\{0.40,0.50,1.70,1.80\}$ 1000
HumanoidStandup $m\in\{0.80,0.90,1.0,1.15,1.25\}$ $d\in\{0.80,0.90,1.0,1.15,1.25\}$ $m\in\{0.60,0.70,1.50,1.60\}$ $d\in\{0.60,0.70,1.50,1.60\}$ $m\in\{0.40,0.50,1.70,1.80,4.00\}$ $d\in\{0.40,0.50,1.70,1.80,4.00\}$ 1000
Walker $m\in\{0.75,1.0,1.25\}$ $d\in\{0.75,1.0,1.25\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ $m\in\{0.20,0.40,1.60,1.80,4.00\}$ $d\in\{0.20,0.40,1.60,1.80,4.00\}$ 2000
Crippled Walker $m\in\{0.75,1.0,1.25\}$ $d\in\{0.75,1.0,1.25\}$ Crippled Joints (right leg) =$=\{0,1,2\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ Crippled Joints (right leg) $\in$ $\{0,1,2\}$ $m\in\{0.20,0.40,1.60,1.80,4.00\}$ $d\in\{0.20,0.40,1.60,1.80,4.00\}$ Crippled Joints (left leg) $\in$ $\{3,4,5\}$ 2000
Crippled Ant $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ Crippled Legs $R_{1}=\varnothing$ or $R_{1}\in\{0,1,2,3\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ Crippled Legs $R_{1}\in\{0,1,2,3\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ Crippled Legs $\{R_{1},R_{2}\}\in\{0,1,2,3,4,5\}$ $(R_{1}\neq R_{2})$ 2000
Crippled Half-cheetah $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ Crippled Joints (front leg) $R_{1}\in$$\{3,4,5\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ Crippled Joints (back leg) $R_{1}\in\{0,1,2\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ Crippled Joints $\{R_{1},R_{2}\}\in\{0,1,2,3,4,5\}$ $(R_{1}\neq R_{2})$ 2000